OpenStudy (anonymous):

Simplify log2x + log3x2

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

So you want to simplify: \[\log(2x) + \log(3x ^{2})\]

OpenStudy (anonymous):

Do you remember any of the rules for logarithms?

OpenStudy (anonymous):

Like how: log(ab) = log(a) + log(b) ?

OpenStudy (anonymous):

Which also implies that: \[\log(x ^{n}) = n*\log(x)\]

OpenStudy (anonymous):

But do you know what "x" is equal to?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

You said it was: \[\log(2x) + \log(3x^{2}) \] right? Well, you CAN simplify that expression. But, unless you know what "x" is equal to (or, what the entire expression is equal to) you will still have the "x" term somewhere in your result.

OpenStudy (anonymous):

It's for a multiple choice question and they all have an X in the answer. But I wanna now how to solve the question rather than just have the answer.

OpenStudy (anonymous):

first, look at that first rule I mentioned (i.e. that log(ab) = log(a) + log(b) ).

OpenStudy (anonymous):

That means that when you see something like: log(2x) You can immediately break that down into: log(2) + log(x)

OpenStudy (anonymous):

would it be log5X^2

OpenStudy (anonymous):

Quick question: Do the multiple choice questions look like "log(a) + b*log(x)" or is there just one "log"?

OpenStudy (anonymous):

there are two logs

OpenStudy (anonymous):

\[\log2x + \log3x ^{2}\]

OpenStudy (anonymous):

Sorry, I meant the answers--do they feature two logs or one?

OpenStudy (anonymous):

no, just one.

OpenStudy (anonymous):

First, because: log(a) + log(b) = log(a*b) That means that: log(2x) + log(3x^2) = \[\log( 2x* 3x ^{2})\]

OpenStudy (anonymous):

You with me so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

log6\[\log6x ^{2}\]

OpenStudy (anonymous):

Wait, you're forgetting one of the "x"s

OpenStudy (anonymous):

Because you're multiplying "2x" and "3x^2", right?

OpenStudy (anonymous):

oh okay, so add the other x to the power? \[\log6x ^{3}\]

OpenStudy (anonymous):

Yeah, you got it. Because 2x * 3x^2= \[6x ^{3}\]

OpenStudy (anonymous):

Also, as an aside, you can also rewrite it as: \[\log(6x ^{3}) = \log(6) + 3*\log(x)\]

OpenStudy (anonymous):

(Sometimes in math it's more helpful to separate the constants from the variables as well as to get rid of the exponents)

OpenStudy (anonymous):

Makes more sense now. Thanks for help

OpenStudy (anonymous):

Ok, glad it's making sense. Good luck!